Trionfale Sightseeing Company (TSC), located in San Marino, owns a large fleet of small airplanes to provide transportation services to the executives of the area corporations. The company is now considering to purchase of a new small airplane for $320,000 to replace the old one. In addition, the company needs to pay an additional $30,000 to install an advanced anti-icing system in the airplane for safety purposes. The company’s old small airplane has a book value of $85,000, but can be sold for $98,000. It is going to be depreciated at a rate of $13,500 per year for four more years to zero salvage vale under the old depreciation method.

The new small airplane will be depreciated using the 5-year MACRS schedule (please use the table below) and it would increase TSC’s before-tax revenues by $71,000 per year but would also increase operating costs by $10,000 per year due to the more sophisticated maintenance. The new small airplane will be sold after 6 year for $210,000. The purchase of the small airplane will require an increase in net operating working capital of $20,000. The net operating working capital will be recovered after the new small airplane is sold. TSC is in the 34% tax bracket and TSC’s weighted average cost of capital is 12 percent. What is NPV of this project, should TSC accept it? You must show your work and explain the approach that you take to find the solution (6 lines). Work without explanation receives zero.
Answer #1
Book Value of old airplane $85,000
Salvage value (brefore tax) $98,000
Gain on salvage =98000-85000 $13,000
Tax on gain (34%*13000) $4,420
Net Cash inflow on salvage=98000-4420= $93,580
a Initial Cash Outlay:
Cost of new airplane $320,000
installation of safety system $30,000
Total Depreciable asset $350,000
Net Salvage cashflow of airplane now ($93,580)
Initial Outlay:350000-93580 $256,420
b DEPRECIATION OF OLD MACHINE
Year from today 1 2 3 4
D1 Annual Depreciation expense $13,500 $13,500 $13,500 $13,500 (NOTE: There is some error in the question.$13500*4=$54000)
(Book Value mentioned $85000. Also mentioned that it is depreciated to zero in 4 years
(We are ignoring this anomaly, in absence of futher information)
DEPRECIATION OF NEW MACHINE
Year(from today) 1 2 3 4 5 6
A Depreciation Rate 20.00% 32.00% 19.20% 11.52% 11.52% 5.76%
D2=A*$350000 Depreciation amount $70,000 $112,000 $67,200 $40,320 $40,320 $20,160
D=D2-D1 Change in annual depreciation expense $56,500 $98,500 $53,700 $26,820 $40,320 $20,160
Salvage Value of new airplane(Before Tax) $210,000 Present value of Cash Flow- (Cash Flow)/((1+i)^N) j=discount Rate -WACC 12 %=0.12 N-Year of Cash Flow C Year From Today 2 E D
Gain on Salvage $210,000
Tax on Gain =210000*34% $71,400
After tax Salvage cash flow=210000-71400 $138,600
Salvage Value of old machine at end of 6 years $0
Incremental Salvage cash flow $138,600